OpenStudy (anonymous):

PROBLEM SOLVING USING SYSTEMS OF EQUATIONS The length of a rectangle is 5cm more than twice the width. The perimeter of the rectangle is 34 cm. Find the length and the width of the rectangle.

OpenStudy (cathyangs):

L=5+2w 2L+2w=34 substitute for L

OpenStudy (lgbasallote):

let L = length of rectangle W = width of rectangle P = perimeter L = 5 + 2w 2L + 2W = 34 L - 2W = 5 2L +2W = 34 3L = 39 L = 13 W = 4

OpenStudy (anonymous):

how would you do that with x and y?

OpenStudy (cathyangs):

just use x and y instead of L and W

OpenStudy (anonymous):

thankyou!